Refine Assumptions for PPE Forecast
Provided below is FY2016 information for Medtronic PLC.

Medtronic plc
Consolidated Statement of Income
($ millions) Apr. 29, 2016
Net sales $29,610
Costs and expenses
Cost of products sold 9,142
Research and development expenses 2,224
Selling, general, and administrative expense 9,469
Special charges (gains), net 70
Restructuring charge, net 290
Certain litigation charges, net 26
Acquisition-related items 283
Amortization of intangiable assets 1,931
Other expense, net 107
Operating profit 6,068
Interest expense, net 955
Income from operations before income taxes 5,113
Provision for income taxes 971
Net income $4,142


Medtronic plc
Consolidated Balance Sheets
($ millions) Apr. 29, 2016 Apr. 24, 2015
Current assets
Cash and cash equivalents $3,098 $5,065
Investments 9,758 14,637
Accounts receivable 5,562 5,112
Inventories 3,473 3,463
Tax assets 697 1,335
Prepaid expenses and other current assets 1,234 1,454
Total current assets 23,822 31,066
Property, plant, and equipment, net 5,063 4,921
Goodwill 41,500 40,530
Other intangible assets, net 26,899 28,101
Long-term tax assets 1,383 774
Other assets 1,559 1,737
Total assets $100,226 $107,129
Current liabilities
Short-term borrowings $1,215 $2,656
Accounts payable 1,709 1,610
Accrued compensation 1,712 1,611
Accrued income taxes 566 935
Deferred tax liabilities - 119
Other accrued expenses 2,185 2,464
Total current liabilities 7,387 9,395
Long-term debt 30,247 33,752
Long-term accrued compensation 1,759 1,535
Long-term accrued income taxes 2,903 2,476
Long-term deferred tax liabilities 3,729 4,700
Other long-term liabilities 1,916 1,819
Total liabilities 47,941 53,677
Shareholders' equity
Ordinary shares - -
Retained earnings 54,153 54,636
Accumulated other comprehensive (loss) (1,868) (1,184)
Total shareholders' equity 52,285 53,452
Total liabilities and shareholders' equity $100,226 $107,129



a. Use the financial statements along with the additional information below to forecast property, plant and equipment, net for FY2017.

CAPEX in FY2016 $1,112 million
Depreciation expense in FY2016 956 million
Forecasted FY2017 net sales 35,953 million



Round to the nearest million.

Forecasted PPE, net for FY2017 $Answer

million

b. Suppose the company discloses in a press release that accompanies its year-end SEC filing that anticipated CAPEX for FY2017 is $1.5 billion. Use the guidance to refine your forecast of property, plant and equipment, net for FY2017.

$Answer

million
Answer #1

Answort @ property, plant & Equipment to of net sale - $5,062 x 100 $ 24,610 =17711 property, plant and equipment, net 2017 =@ Revised Depreciation expense for 2017 -(5063 +1500 1x164 = 1050 Revised forecasted PPE, net for 2017 = 5063 +1500-1050 =$55